### v1

# Observational Auditing of Label Privacy

Abstract
Differential privacy (DP) auditing is essential for evaluating privacy guarantees in machine learning systems. Existing auditing methods, however, pose a significant challenge for large-scale systems since they require modifying the training dataset—for instance, by injecting out-of-distribution canaries or removing samples from training. Such interventions on the training data pipeline are resource-intensive and involve considerable engineering overhead. We introduce a novel observational auditing framework that leverages the inherent randomness of data distributions, enabling privacy evaluation without altering the original dataset. Our approach extends privacy auditing beyond traditional membership inference to protected attributes, with labels as a special case, addressing a key gap in existing techniques. We provide theoretical foundations for our method and perform experiments on Criteo and CIFAR-10 datasets that demonstrate its effectiveness in auditing label privacy guarantees. This work opens new avenues for practical privacy auditing in large-scale production environments.
[…]
Iden Kalemaj
1 Introduction
Differential privacy (DP) auditing has become an important tool for evaluating privacy guarantees in machine learning systems. Recent advances in auditing methods that require only a single run have made it feasible to evaluate privacy for large-scale models without prohibitive computational costs (Steinke et al., 2024; Mahloujifar et al., 2025b). However, most existing auditing approaches still require modifying the training dataset by injecting known entropy or canary data, which limits their applicability in industry-scale environments, where modifications to the training data pipeline require significant engineering overhead.
In this work, we propose a novel auditing methodology that eliminates the need for dataset modification. Our approach enables privacy evaluation using the natural nondeterminism present in the data distribution itself. We formalize and empirically validate this methodology in the setting of auditing Label DP, generalizable to privacy guarantees for any protected attribute. This capability addresses a significant gap in current auditing techniques, which have primarily focused on membership inference attacks (Shokri et al., 2017; Carlini et al., 2022) rather than attribute inference. In particular, existing methods for auditing Label DP either require adding out-of-distribution canaries to the training set (Malek et al., 2021) or are applicable only to a limited set of mechanisms (Busa-Fekete et al., 2024).
Observational privacy auditing cannot be done unconditionally, without making certain assumptions about the underlying distribution (Hernán and Robins, 2020, Chapter 3). Unlike anecdotal instances of privacy violations (Barbaro and Zeller Jr, 2006; Narayanan and Shmatikov, 2008; Carlini et al., 2021), auditing seeks to provide statistically valid measurements of memorization. In other words, the objective of privacy auditing is to establish causality—demonstrating that a model behaves in a certain way because it was trained on specific data. Traditionally, most rigorous membership inference attacks establish and measure causal effects through randomized control trials (RCTs), which require interference with the training data. By reframing auditing as a security game between two adversarial parties, our approach eliminates the need for training-time intervention.
Our key assumption is the availability of a distribution that approximates the ground truth. Concretely, the Label DP auditing mechanism relies on access to a proxy label-generating distribution. The proxy does not need to match the ground truth distribution, provided the adversary cannot distinguish between them (with reasonable computational resources). Under this assumption, the counterfactual examples generated by the challenger can be used to evaluate the attack on the model’s claimed Label DP guarantees without training data intervention. The attack can be made practical by using a model other than the target model as the proxy distribution. Further, in an incremental learning setting, earlier model checkpoints can be used as the proxy distribution, thus requiring no additional model training and minimal engineering overhead. The
We demonstrate that our observational auditing framework provides accurate privacy bounds that match those obtained by interventional methods and provable privacy guarantees. Through a series of cryptography-inspired games we establish the theoretical foundations for auditing privacy without training data manipulation. By lowering the complexity of privacy auditing, our approach enables its application in a wider variety of contexts.
2 Background and prior work
Differential Privacy (DP) introduced by Dwork et al. (2006) is a leading framework for providing rigorous privacy guarantees in statistical data analysis and machine learning. In its standard formulation, DP bounds the impact that any single individual’s record has on the outcome of a computation by constraining how much the distribution of outputs may differ between neighboring datasets. The definition’s strong theoretical guarantees, resilience to arbitrary auxiliary data, and compositional properties have driven its adoption in academic research and industry deployments (Fioretto and Hentenryck, 2025).
The general DP framework can be adapted to settings where only certain parts of a dataset are considered private. This paper focuses on Label DP, which has emerged as an important objective for PPML, particularly in the domain of recommendation systems (Chaudhuri and Hsu, 2011; Ghazi et al., 2021; Malek et al., 2021; Wu et al., 2023). The following factors motivate Label DP as a uniquely valuable privacy concept:
  • The label—representing the user’s choice, expressed preference, or the outcome of an action—may be the only sensitive part of the record, with the rest being publicly available, static, non-sensitive data.
  • In machine learning, labels are particularly vulnerable to memorization compared to other attributes since they most directly influence the loss function.
  • In settings with mixed public/private features, instead of applying privacy-preserving techniques to sensitive features, one may exclude them from the model, potentially sacrificing some accuracy. In supervised learning, however, labels are indispensable—there is no analogous alternative to omitting sensitive features. In further separation, once training completes, labels can be safely discarded, whereas features must be available for inference.
Complementing the strong worst-case guarantees of DP that bound the privacy loss from above on all inputs, privacy auditing empirically measures the privacy loss on concrete instances, providing a lower bound on DP’s numerical parameters. Privacy auditing can be used for finding bugs in claimed implementations of DP algorithms (Ding et al., 2018), advancing understanding of complex DP mechanisms (Malek et al., 2021; Nasr et al., 2023, 2025), or guardrailing models in a production environment (Agrawal and Book, 2025).
Privacy auditing consists of two components: a privacy game between the challenger and the attacker, and an auditing analysis that translates the attacker’s success into lower bounds on the $({\epsilon},\delta)$-DP guarantee (or other forms of DP). The privacy game is characterized by the capabilities and resources of the parties, and the attacker’s goals, such as reconstruction, membership or attribute inference. Auditing analyses, typically used for membership inference, can be applied to any stochastic privacy game (Swanberg et al., 2025). We show theoretical results for our label inference attack building on Steinke et al. (2024) and Mahloujifar et al. (2025b).
Membership inference attacks (MIAs)—where an adversary uses model access and knowledge of the data distribution to determine whether a sample was part of training—have received significant attention in the literature (Shokri et al., 2017; Yeom et al., 2018; Salem et al., 2019; Sablayrolles et al., 2019; Song and Shmatikov, 2019; Nasr et al., 2019; Leino and Fredrikson, 2020; Carlini et al., 2022; Ye et al., 2022; Zarifzadeh et al., 2024; Bertran et al., 2023). This attack category directly maps to the differential privacy guarantee where two neighboring datasets differ in the presence of one training sample.
A less studied category is attribute inference attacks (Yeom et al., 2018) where the adversary reconstructs a protected attribute given access to a partial record, of which label inference attacks are a special case (Malek et al., 2021; Busa-Fekete et al., 2024). A difficulty and common pitfall with such attacks is to properly account for the adversary’s baseline success, achieved by exploiting knowledge of the data distribution and correlations between the public and protected attributes (Jayaraman and Evans, 2022). The label inference attack of Malek et al. (2021) uses canaries with random binary labels, which sets the adversary’s baseline accuracy to $0.5$ and allows Label DP to be audited via standard MIA analyses. However, this attack interferes with the training dataset and can affect model performance if too many out-of-distribution samples are injected.
Most auditing methods in the literature are “interventional,” as their privacy game involves manipulating the training dataset: MIAs require excluding a subset of the data from training (Steinke et al., 2024; Mahloujifar et al., 2025b), whereas Malek et al. (2021) modifies the training labels. These requirements severely restrict the applications of these auditing methods. Instead, our label inference attack can run entirely post-training. MIAs can be stated as observational privacy games if the challenger is able to sample fresh samples from the distribution (Ye et al., 2022). However, obtaining new samples from the distribution (or from its close approximation), without affecting the training pipeline, remains an extremely challenging open problem (Meeus et al., 2025).
Kazmi et al. (2024) propose an observational MIA that uses generative models such as generative adversarial networks (GANs) to obtain fresh samples. Differently from Kazmi et al. (2024), we propose a more general observational auditing framework (with observational MIA and LIA as special cases), and provide theoretical lower bounds on the privacy guarantees for these general attacks. A challenge with using synthetic samples is to account for how the shift between the generative model and the true distribution boosts the success of the attacker. Counterfactual labels are much easier to generate than entire samples (models are trained to do precisely this), and we can assume that an adversary with limited computation resources cannot distinguish between a label from the distribution and a label predicted from the best model the adversary can train.
The closest observational label inference attack to ours is the recent work of Busa-Fekete et al. (2024), which measures the label reconstruction advantage of the adversary with and without access to the model. This metric is not translated into a lower bound on $\epsilon$} in the differential privacy guarantee; in fact, such translation would be difficult because the adversary has a different baseline success (prior) for each sample. Additionally, this approach requires estimating probabilities of the mechanism’s output given a particular label, limiting its applicability to simple mechanisms like randomized response and random label aggregation. In contrast, our label inference attack is observational and applicable to all mechanisms. It can audit label privacy in a statistically valid manner because it sets up a game where the baseline accuracy of the adversary is $0.5$ for all samples.
3 Preliminaries
Notation
We use calligraphic letters such as $\mathcal{X},\mathcal{Y},\mathcal{D}$ to denote sets and distributions. Capital letters such as $X,Y,D$ denote random variables and datasets, and lowercase letters denote their values. $\mathcal{X}^{n}$ is the set of all datasets of size $n$ with elements from $\mathcal{X}$, whereas $\mathcal{X}^{*}$ is the set of all finite-size data sets with elements from $\mathcal{X}$.
In this work, we audit the simulation-based definition of differential privacy. It generalizes the traditional add/remove (or “leave-one-out”, or “zero-out”) notion of DP to support a privacy unit that is a subset of the sample’s attributes. The definition compares the distribution of a mechanism $M$ on dataset $D$ with that of a simulator that emulates the output of $M$ on $D$ without seeing the protected attributes of a record. See Appendix 8 for further discussion of this definition.
Definition 1 (Simulation-based privacy for protected attributes (Mahloujifar et al., 2025a)).
Let records $(x,y)\in\mathcal{X}\times\mathcal{Y}$ be such that $x$ is public or non-sensitive and thus need not be protected. We say that a randomized mechanism $M\colon(\mathcal{X}\times\mathcal{Y})^{*}\to\mathcal{Z}$ is $(\epsilon,\delta)$-SIM-DP with respect to a simulator $\mathrm{Sim}\colon(\mathcal{X}\times\mathcal{Y})^{*}\times\mathcal{X}\to\mathcal{Z}$ if for all datasets $D\in(\mathcal{X}\times\mathcal{Y})^{*}$, $(x,y)\in D$, and $D^{\prime}=D\setminus\{(x,y)\}$, it holds
$M(D)\approx_{{\epsilon},\delta}\mathrm{Sim}(D^{\prime},x).$ (1)
For the more advanced notion of $f$-DP (Dong et al., 2020), a mechanism is f-SIM-DP if
$M(D)\approx_{f}\mathrm{Sim}(D^{\prime},x).$ (2)
We also say $M$ is $(\epsilon,\delta)$-SIM-DP (resp. f-SIM-DP) if there exists a simulator $\mathrm{Sim}$ for which (1) (resp. (2)) holds.
Our auditing guarantees are stated for a family of generic simulators that treat $M$ as a black-box. Such a simulator imputes the missing part $y$ of the record based on the public part $x$ and runs the original mechanism.
Definition 2 (Imputation-based simulator).
For a mechanism $M\colon(\mathcal{X}\times\mathcal{Y})^{*}\to{\mathcal{Z}}$, a data distribution $\mathcal{D}$ supported on $\mathcal{X}\times\mathcal{Y}$, a dataset $D\in(\mathcal{X}\times\mathcal{Y})^{*}$ and public part of a record $x\in\mathcal{X}$, the imputation-based simulator $\mathrm{Sim}_{M,\mathcal{D}}$ is defined as
$\mathrm{Sim}_{M,\mathcal{D}}(D,x)\mathrel{\triangleq}M(D\cup\{(x,y^{\prime}\}),\quad{\text{where }}y^{\prime}\sim\mathcal{D}\mid x.$
Finally, we define empirical privacy auditing. A similar definition holds for $(\epsilon,\delta)$-SIM-DP.
Definition 3 (Auditing simulation-based DP).
An audit procedure takes the description of a mechanism $M$, a trade-off function $f$, a simulator $\mathrm{Sim}$ and decides whether the mechanism satisfies f-SIM-DP. We define it as a two-step process.
  • $\mathrm{Game}\colon\mathcal{M}\times\mathcal{S}\to\mathcal{O}$. The auditor runs a potentially randomized experiment/game using the description of mechanism $M\in\mathcal{M}$ and the simulator $\mathrm{Sim}$. The auditor receives the game output $o\in\mathcal{O}$.
  • $\mathrm{Evaluate}\colon\mathcal{O}\times\mathcal{F}\to\{0,1\}$. The output is 0 if the auditor rejects the hypothesis that $M$ satisfies f-SIM-DP based on evidence $o$, and $1$ otherwise.
4 Observational versus interventional privacy games
An observational privacy game considers the training dataset as a given. In contrast, an interventional privacy game interferes with the training data pipeline and the eventual dataset used for training. In this section, we formalize our observational privacy auditing framework. To that end, we introduce a generic attack game, generalizing Swanberg et al. (2025), as Algorithm 1.
The privacy game occurs between two parties: a challenger and an adversary. A key difference from Swanberg et al. (2025) is that we split the challenger algorithm into two stages: determining the training data $(G_{\text{intervention}})$ and determining the additional outputs provided to the adversary ($G_{\text{hint}}$). With this, we can separate observational games from interventional ones. Further we distinguish between the training dataset $D$ and additional game artifacts $S$, which are used by the challenger to set up a stochastic game (e.g., sampling random bits).
Algorithm 1 Generic Attack Game (adapted from Swanberg et al. (2025))
0: Mechanism $M(\cdot)$, data distribution $\mathcal{D}$, distribution for game artifacts $\mathcal{D}_{\mathsf{prior}}$, adversary AA
1: Sample training dataset $D=(x_{1},\dots,x_{m})$ where $x_{i}\sim\mathcal{D}$.
2: Sample game artifacts $S\sim\mathcal{D}_{\mathsf{prior}}$.
[…]
4: Let $o_{1}\leftarrow M(\overline{D})$.
[…]
6: Run attack $A(o_{1},o_{2})$ with access to $\mathcal{D}$, $\mathcal{D}_{\mathtt{prior}},M$.
7: Measure adversary success with loss metric $\mathcal{L}(A(o_{1},o_{2}),S)$.
The algorithm $G_{\text{intervention}}$ determines the training dataset for $M$, obtained from a potential modification of the fixed training set $D$. For instance, in the one-run MIA (Steinke et al., 2024) the artifacts are $S=(S_{i}\sim\mathsf{Bernoulli}(0.5)\colon i\in[m])$, where $m$ is the number of canaries. The training dataset $\overline{D}$ is obtained from $D$ by including all samples $x_{i}\in D$ where $S_{i}=1$ and excluding all samples where $S_{i}=0$.
The role of $G_{\text{hint}}$ is to collect additional information the challenger provides to the adversary, based on game artifacts, training data, and the output of the trained model $M(\overline{D})$. In the one-run MIA, $o_{2}$ is the vector of targets $x_{1},\dots,x_{m}$ from the training set $D$.
Definition 4 (Observational games).
We call a privacy game, as outlined in Algorithm 1, observational if $\overline{D}=D$ (and as a result $o_{1}=M(D)$). That is, $M$ is trained on the original $D$, and the observations of the adversary consist of (1) output of $M(D)$ and (2) additional postprocessing of $D$, $M(D)$ according to the game artifacts $S$.
Observational versus interventional MIA. We have described how the (interventional) one-run MIA (Steinke et al. (2024)) can be framed as Algorithm 1. We now present an observational one-run MIA (following Ye et al. (2022)).
Given data distribution $\mathcal{D}$, sample a sequence $S$ of game artifacts $S_{i}=(b_{i},x_{i}^{\prime})\sim\{0,1\}\times\mathcal{D}$ for $i\in[m]$. The bits $b_{i}\sim\mathsf{Bernoulli}(0.5)$, whereas $x_{i}$ is a fresh sample from the distribution (which is highly unlikely to be in $D)$. Then, train model $M$ on $\overline{D}=D$. Let $o_{2}[i]=x_{i}$ if $b_{i}=0$ and $o_{2}[i]=x_{i}^{\prime}$ if $b_{i}=1$ for $i\in[m]$. That is, the adversary receives either a training sample $x_{i}$ or a sample $x_{i}^{\prime}$ from the distribution with probability $0.5$. The adversary has to guess $b_{i}$, i.e., which of the samples it is seeing. This game is observational because the training dataset for $M$ remains unchanged.
From the adversary’s perspective, the observational one-run MIA has the same distribution as the interventional MIA. For the challenger, the difference matters as the observational game does not alter the training pipeline. In this game, the source of counterfactual samples $x_{i}^{\prime}$ can be a distribution that approximates $\mathcal{D}$ sufficiently well, as is the case with the attack of Kazmi et al. (2024) who use generative models to sample $x_{i}^{\prime}$.
5 Observational attribute inference
In this section, we describe our observational attribute inference attack. It allows privacy measurement with respect to any set of protected attributes, which can be the entire record (as in the observational MIA, Section 4) or just the label, for Label DP auditing. We provide theoretical results for obtaining empirical privacy lower bounds from our game. In particular, our analysis provides lower bounds on simulation-based DP in the add/remove privacy model (see Appendix 8).
Algorithm 2 Observational attribute inference in one run
0: Oracle access to a mechanism $M(\cdot)$, data distribution $\mathcal{D}$ and approximate distribution $\mathcal{D}^{\prime}$ supported on $\mathcal{X}\times\mathcal{Y}$, attacker AA
1: Let $D^{0}=\Big((x_{1},y_{1}^{0}),\dots,(x_{m},y_{m}^{0})\Big)$, where $(x_{i},y_{i}^{0})\sim\mathcal{D}$ for $i\in[m]$.
2: Run mechanism $M$ on $D^{0}$ to get output $o_{1}$.
3: Sample game artifacts $\Big((b_{1},y_{1}^{1}),\dots,(b_{m},y_{m}^{1})\Big)$ such that $(b_{i},y_{i}^{1})\sim\mathsf{Bernoulli}(0.5)\times\mathcal{D}^{\prime}\mid x_{i}$.
4: Construct a dataset $D^{b}=\Big((x_{1},y_{1}^{b_{1}}),\dots,(x_{m},y_{m}^{b_{m}})\Big).$
5: Run attack $A$ with input $o_{1}=M(D^{0})$, $o_{2}=D^{b}$, and access to $\mathcal{D}$, $\mathcal{D}^{\prime}$.
6: Reconstruct a vector of predictions $b^{\prime}=(b^{\prime}_{1},\dots,b^{\prime}_{m})$ which is supported on $\{0,1,\bot\}^{m}$.
7: Count $c$, the number of correct guesses where $b_{i}^{\prime}=b_{i}$, and $c^{\prime}$, the total number of guesses where $b_{i}^{\prime}\neq\bot$.   $\triangleright\bot$ indicates abstention from guessing
8:return $(c,c^{\prime})$.
Similar to prior auditing papers (Mahloujifar et al., 2025b; Steinke et al., 2024) the adversary can choose to abstain from guessing on samples where it is least confident, to boost its positive likelihood ratio. The observational game can use the entire dataset as canaries (i.e., $m=n$).
Obtaining approximate distributions. A key aspect in implementing the observational attribute inference attack is to produce the proxy distribution $\mathcal{D}^{\prime}$ from which the counterfactual partial records $y_{i}^{1}$ are sampled. One option is to train an additional model $M^{\prime}$ to predict the missing attribute(s). For label inference attacks, which are a special case of Algorithm 2, we sample counterfactual label $y_{i}^{i}$ from $\mathsf{Multinoulli}(M^{\prime}(x_{i}))$. In an online machine learning system, where the model trains incrementally as more recent data becomes available, one can use a prior model checkpoint as the model $M^{\prime}$ (and run the attack on the newer data). This eliminates the need for training any additional models, making our proposed label inference attacks very lightweight in terms of computational overhead and implementation complexity.
Auditing guarantees when $\mathcal{D}=\mathcal{D}^{\prime}$. We first establish auditing guarantees for the simpler case where $\mathcal{D}=\mathcal{D}^{\prime}$, i.e., when we can sample counterfactual records from the ground truth distribution. Our bounds are stated for auditing $f$-DP, following an analogous argument in Mahloujifar et al. (2025b), which gives the tightest DP auditing analysis and can be translated to the language of $({\epsilon},\delta)$. The auditing applies an evaluation procedure, Algorithm 3, described and analyzed in Appendix 9.
Theorem 5 (Auditing $f$-DP with no distribution shift).
Let $M\colon(\mathcal{X}\times\mathcal{Y})^{*}\to\mathcal{Z}$ be a mechanism, $\mathcal{D}$ the data distribution, and $\mathrm{Sim}_{M,\mathcal{D}}$ the imputation-based simulator (Definition 2). Let $C=\sum_{i\in[m]}\mathbf{1}[b_{i}^{\prime}=b_{i}]$ be the total number of correct answers from the one-run observational attribute inference attack with no distribution shift (Algorithm 2, when $\mathcal{D}^{\prime}=\mathcal{D}$) for an adversary that makes $c^{\prime}$ guesses. If $M$ is f-SIM-DP and Algorithm 3 returns True on $(c^{\prime},c,M,f,\gamma=0.05)$, then $\Pr[C\geq c]\leq\gamma.$
Auditing guarantees under distribution shift. Now we consider the more general case when $\mathcal{D}\neq\mathcal{D}^{\prime}$. We state an information-theoretic bound, which depends on the total variation (TV) distance between $\mathcal{D}$ and $\mathcal{D}^{\prime}$. Intuitively, the larger the distance between the two distributions, the weaker the lower bound we can obtain on the privacy guarantee. The value $\tau$ in Theorem 6 is an upper bound on the adversary’s a priori (i.e., before having access to the target model) success probability in distinguishing whether a sample $(x,y^{b})$ is from $\mathcal{D}$ or $\mathcal{D}^{\prime}$.
Theorem 6 can also be stated in terms of the adversary’s ability to distinguish between $\mathcal{D}$ and $\mathcal{D}^{\prime}$ given its resource constraints. Such a bound is particularly meaningful for Label DP auditing, where we can assume that the adversary cannot distinguish $y\mid x$ from $\mathsf{Multinoulli}(M^{\prime}(x))$ if $M^{\prime}$ is the best classifier on $\mathcal{D}$ available to the adversary.
We state here a bound akin to Steinke et al. (2024), that is easier to interpret in the case of distribution shift. We present a bound on auditing $f$-DP under distribution shift in Appendix 9.
Theorem 6 (Auditing $({\epsilon},\delta)$-DP under distribution shift).
Let $M\colon(\mathcal{X}\times\mathcal{Y})^{*}\to\mathcal{Z}$ be a mechanism, $\mathcal{D}$ the data distribution, $\mathcal{D}^{\prime}$ the approximate distribution, and $\mathrm{Sim}_{M,\mathcal{D}^{\prime}}$ the imputation-based simulator. Let $C=\sum_{i\in[m]}\mathbf{1}[b_{i}^{\prime}=b_{i}]$ be the total number of correct answers from the one-run observational attribute inference attack (Algorithm 2). If $\mathcal{M}$ is $({\epsilon},\delta)$-$\mathrm{Sim}_{M,\mathcal{D}^{\prime}}$-DP and $\mathrm{TV}(\mathcal{D}|x_{i},\mathcal{D}^{\prime}|x_{i})\leq\tau$ for all $x_{i}\in D$, then
$\Pr_{S,D}[C\geq c\mid A(o_{1},o_{2})=b^{\prime}]\leq\Pr_{\hat{C}_{i}\sim\mathsf{Bernoulli}(\beta)}\Big[\sum_{i\in[m]}\hat{C}_{i}\geq c\Big]+n\delta$ (3)
where $\beta=\frac{e^{\epsilon}}{e^{\epsilon}+\frac{1-\tau}{1+\tau}}$.
A full theorem statement and its proof are deferred to Appendix 10.
6 Experiments on auditing Label DP
We validate our theoretical framework with experiments on two real representative datasets: CIFAR-10 (Krizhevsky, 2009) for image classification and Criteo (Criteo AI Lab, 2015) for tabular data with sensitive labels. For each dataset, we train classifiers using standard Label DP mechanisms and empirically evaluate the privacy guarantees using our audit procedure. We also evaluate our Label DP auditing technique for Randomized Response (Warner, 1965) on synthetic data. Our attack is implemented in the PrivacyGuard library (Melis et al., 2025). [^1] https://github.com/facebookresearch/PrivacyGuard
6.1 Label DP learning algorithms
The earliest approach to achieving Label DP is the Randomized Response (RR) mechanism (Warner, 1965). In RR, each training label is randomly replaced according to a fixed probability distribution before being shared with the learning algorithm. This randomization helps protect the privacy of individual labels. We briefly review several recent Label DP mechanisms that improve on RR.
Label Private One-Stage Training (LP-1ST, Ghazi et al. (2021)) Instead of using a fixed distribution as in RR, LP-1ST samples each training label $y_{i}$ from a learned prior distribution $P(y|X_{i})$. The prior can be estimated by observing the top-$k$ predictions from a pretrained model (either in domain or out-of-domain), restricting RR to the most probable labels. Alternatively, the training can be split into multiple stages, where an earlier model provides the prior for the next (LP-MST).
Private Aggregation of Teacher Ensembles with FixMatch (PATE-FM, Malek et al. (2021)) PATE-FM combines the FixMatch semi-supervised learning algorithm (Sohn et al., 2020) with private aggregation. Multiple teacher models are trained, each using all unlabeled examples and disjoint subsets of the labeled data. The predictions from these teachers are then aggregated in a differentially private manner using the PATE framework (Papernot et al., 2017) to train a student model.
Additive Laplace with Iterative Bayesian Inference (ALIBI, Malek et al. (2021)) ALIBI achieves Label DP by adding Laplace noise to the one-hot encoded labels (Ghosh et al., 2012), making the released labels differentially private. The model is trained on soft labels computed from noisy observations via Bayesian inference.
6.2 Attack implementation and adversarial strategy
A key ingredient in implementing our label inference attack is generating the reconstructed label $y^{1}$ given features $x$ (see Algorithm 2). We generate $y^{1}$ from the predictions of a reference model $M^{\prime}$, trained on separate data (but from approximately the same distribution) as the target model $M$. Specifically, $y^{1}\sim\mathsf{Multinoulli}(M^{\prime}(x))$, where $M^{\prime}(x)$ are the predictions of $M^{\prime}$ on $x$ for each class.
For the Criteo dataset, where data is collected over 28 consecutive days, we train $M^{\prime}$ on Day 0 data and the target model on Day 1 data. While there may be some distributional shift between Day 0 and Day 1 data, we assume this is small, and that the adversary cannot distinguish between the true labels $y^{0}$ and the reconstructed labels $y^{1}$ (before having access to the model).
For CIFAR-10 experiments we randomly split the training data (n=50n=50K samples) into two. We train $M^{\prime}$ on the first half, and the target model $M$ on the second half.
We run the attack on m=200m=200K canaries for Criteo and m=10m=10K canaries for CIFAR-10. For the MIA experiments we use the same number of canaries for the non-members (taken from the test set).
The adversary obtains its guesses by computing per-example scores. Let $x$ be the features and $y^{b}\in\{y^{0},y^{1}\}$ be the label received by the adversary. The adversary computes a score that correlates with whether $y^{b}$ is the reconstructed or the training label. The score consists of two components. The first component $s_{1}(x,y^{b})$ is the difference in probabilities that $y^{b}$ came from the training set versus the reconstructed distribution:
$s_{1}(x,y^{b})$ $=\Pr[y^{0}=y^{b}\mid M(x)]-\Pr[y^{1}=y^{b}\mid M^{\prime}(x)]$
$=M(x)[y^{b}]-M^{\prime}(x)[y^{b}],$
where $M(x)[y^{b}]$ is the prediction of $M$ on $x$ for class $y^{b}$. Since the adversary’s performance is measured at the tails of the score distribution, the adversary prefers to guess on samples where $\Pr[y^{0}\neq y^{1}]$ is high. Thus the second component of the score is defined as
$s_{2}(x,y^{b})=\Pr[y^{0}\neq y^{1}]=1-M^{\prime}(x)[y^{b}].$
The final score combines the two components as $s(x,y^{b})=s_{1}(x,y^{b})\cdot s_{2}(x,y^{b})^{t}$ with a hyperparameter $t\geq 0$ that allows for weighting the two components separately. We use $t=2$, as $t>1$ gives tight lower bounds for RR. The adversary guesses on $c^{\prime}\%$ of samples with highest absolute scores. We sweep $c^{\prime}\in\{1,2,\dots,100\}$ and report the highest $\epsilon$} achieved at 95% confidence, averaged over 100 repetitions of the game (resamplings of counterfactual labels).
6.3 CIFAR-10 experiments
For CIFAR-10, we treat the image classes as sensitive labels. We train standard convolutional neural networks with varying privacy budgets ${\epsilon}\in\{1.0,10.0,\infty\}$ (see model accuracy in Table 4, Appendix 12). We then audit Label DP using our observational game and report results in Table 1.
Table 1: CIFAR-10. Auditing Label DP algorithms under different $\epsilon$} with $\delta=10^{-5}$.
Label DP Algorithm CIFAR-10
$\epsilon$$\infty$ ${\epsilon}=10.0$ ${\epsilon}=1.0$
LP-1ST 2.13 $\pm$ .22 2.02 $\pm$ .33 0.43 $\pm$ .05
LP-1ST (out-of-domain prior) 2.26 $\pm$ .22 1.86 $\pm$ .25 0.90 $\pm$ .07
PATE-FM 2.42 $\pm$ .32 2.22 $\pm$ .24 0.79 $\pm$ .09
ALIBI 2.53 $\pm$ .33 2.18 $\pm$ .27 0.67 $\pm$ .07
6.4 Criteo experiments
The Criteo dataset contains user click-through data with demographic information encoded as 13 numerical features and 26 categorical features. A binary label indicates whether the user clicked on the ad. The distribution of the labels is highly imbalanced, with only 3% of positives (clicks). The overall dataset contains over 4 billion click log data points over a period of 24 days. We followed the same setup in Wu et al. (2023) where 1 million data points are selected for each day. We divide the data into 80% for training, 4% for validation, and 16% for testing. Model performance is evaluated using the log-loss metric on the test set.
We train gradient boosting decision trees with the CatBoost library (Prokhorenkova et al., 2018) with varying privacy budgets ${\epsilon}\in\{0.1,1.0,2,0,4.0,8.0,\infty\}$ and evaluate label privacy using observational privacy auditing (Table 2). Table 5 (Appendix 12) shows model performance under different Label DP algorithms and values $\epsilon$}.
Table 2: Criteo. Auditing Label DP algorithms under different $\epsilon$} with $\delta=10^{-5}$. Similarly to Wu et al. (2023), for LP-1ST (domain prior) at ${\epsilon}\in\{0.1,1,2\}$ the training process did not produce meaningful outcomes.
Label DP Algorithm $\epsilon$$\infty$ ${\epsilon}=8$ ${\epsilon}=4$ ${\epsilon}=2$ ${\epsilon}=1$ ${\epsilon}=0.1$
LP-1ST 1.37 $\pm$ .14 1.29 $\pm$ .12 1.22 $\pm$ .17 0.59 $\pm$ .05 0.34 $\pm$ .02 0.06 $\pm$ .01
LP-1ST (domain prior) 1.47 $\pm$ .22 1.37 $\pm$ .10 1.28 $\pm$ .21
LP-1ST (noise correction) 1.31 $\pm$ .12 1.26 $\pm$ .11 1.04 $\pm$ .16 0.52 $\pm$ .07 0.40 $\pm$ .08 0.06 $\pm$ .01
LP-2ST 1.52 $\pm$ .14 1.46 $\pm$ .11 1.24 $\pm$ .15 0.75 $\pm$ .05 0.61 $\pm$ .07 0.06 $\pm$ .01
PATE 1.60 $\pm$ .15 1.48 $\pm$ .14 1.28 $\pm$ .12 0.71 $\pm$ .07 0.59 $\pm$ .05 0.06 $\pm$ .01
6.5 Comparison with Existing Methods
We compare our observational auditing approach against traditional canary-based methods to demonstrate the effectiveness and practicality of our framework. More specifically, we evaluate against the lightweight difficulty calibration MIA in Watson et al. (2022), where membership scores are adjusted to the difficulty of correctly classifying the target sample. For each canary datapoint, we set the calibrated membership scores as the difference in the loss between the target model and the reference model $M^{\prime}$. Fig. 1 shows how our method achieves similar auditing results when compared to MIA on the CIFAR-10 and Criteo datasets.
(a) CIFAR-10
(b) Criteo
Figure 1: Comparison with MIA for different Label DP Algorithms on CIFAR-10 and Criteo datasets. The error bar represents the standard deviation across 100 different repetitions.
We leave for future work a comparison with more computationally intensive methods, such as Zarifzadeh et al. (2024), which require training multiple auxiliary (shadow) models to achieve state-of-the-art attack performances.
Figure 2: Auditing randomized response when the adversary guesses on 0.1% of samples. The counterfactual labels are generated either from the ground-truth distribution or a proxy distribution obtained from the predictions of logistic regression model.
6.6 Synthetic data and auditing randomized response
We empirically demonstrate the tightness of our Label DP auditing algorithm for Randomized Response using synthetic data. The distribution consists of $k$ balanced classes, each generated from a $5$-dimensional Gaussian with the same covariance but a shifted mean. We generate $n=10^{6}$ samples and experiment with $k\in\{2,5,10\}$. Counterfactual labels $y^{1}$ are generated using either the true distribution or the predictions of a logistic regression model. Figure 2 shows empirical epsilon lower bounds at $95\%$ confidence when the adversary makes $0.1\%$ non-abstaining guesses. We obtain tight lower bounds for ${\epsilon}\in[1,4]$. At lower epsilons the audit overestimates privacy loss due to a higher variance induced by a small number of guesses. Using more guesses at lower epsilon fixes the issue (Appendix 11). Obtaining tight lower bounds at very high epsilon is a limitation of current auditing methodology (Steinke et al., 2024; Mahloujifar et al., 2025b).
7 Discussion
In this paper, we establish a framework for auditing privacy without any intervention during the training process. This enables a principled privacy evaluation in settings where the training process is outside the control of the privacy auditor. This may sound counterintuitive as privacy auditing is a form of causal analysis. However, our method can provide provable guarantees on auditing performance under certain assumptions about the data distribution. We envision that our framework will broaden the scope of privacy auditing applications, as it does not require any supervision of the training process and can be conducted by third parties.